# Taipei 10617, Taiwan

### v2

@@ 4 $\mathcal{E}(\partial_{-}\mathcal{E}(\ell_{n_{1}})\mathcal{E}(\ell_{m}))=0$
To begin, recall that $\mathcal{E}(\partial_{-}\mathcal{E}(\ell_{n_{1}})\mathcal{E}(\ell_{m}))=0$ is a polynomial equation which involves $j(x^{-})$ and its derivatives up to order $2(n_{1}+m)$. So if $\mathcal{E}(\partial_{-}\mathcal{E}(\ell_{n_{1}})\mathcal{E}(\ell_{m}))=0$, then any derivative of $\mathcal{E}(\partial_{-}\mathcal{E}(\ell_{n_{1}})\mathcal{E}(\ell_{m}))$, say with respect to $j_{2(n_{1}+m)-p}$ with $0\leq p\leq 2(n_{1}+m)$, is zero
@@
We first focus on (4.1) when $0\leq p<2n_{1}$. Using the above and the commutation relation $[\frac{\partial}{\partial j_{k}},\partial_{-}]=\frac{\partial}{\partial j_{k-1}}$, (4.1) becomes
@@
$+\sum_{k=0}^{p}\sum_{l=0}^{k}\Bigg{(}{2m-p+1+k\choose k+1-l}\partial_{-}^{k+1-l}\frac{\partial\mathcal{E}(\ell_{n_{1}})}{\partial j_{2n_{1}-l}}\frac{\partial\mathcal{E}(\ell_{m})}{\partial j_{2m-p+k}}-{2n_{1}+1-l\choose k+1-l}\partial_{-}^{k+1-l}\frac{\partial\mathcal{E}(\ell_{m})}{\partial j_{2m-p+k}}\frac{\partial\mathcal{E}(\ell_{n_{1}})}{\partial j_{2n_{1}-l}}\Bigg{)}$
@@
For $0\leq p<2n_{1}$, this is a recursion relation for $\frac{\partial\mathcal{E}(\ell_{m})}{\partial j_{2m-p}}$ of order $p$, which can be used to solve for $\frac{\partial\mathcal{E}(\ell_{m})}{\partial j_{2m-p}}$ in terms of $\frac{\partial\mathcal{E}(\ell_{n_{1}})}{\partial j_{2n_{1}-l}}$ with $0\leq l\leq p$.
We simplify the recursion relation further by expanding out the derivatives $\frac{\partial\mathcal{E}(\ell_{n_{1}})}{\partial j_{2n_{1}-l}}$ and $\frac{\partial\mathcal{E}(\ell_{m})}{\partial j_{2m-l}}$, and noting that they can be written in the following form
$\frac{\partial\mathcal{E}(\ell_{n_{1}})}{\partial j_{2n_{1}-l}}=\sum_{i=0}^{\lfloor l/2\rfloor}{n_{1}-i\choose l-2i}\partial^{l-2i}_{-}A_{n_{1}-i},$ $\frac{\partial\mathcal{E}(\ell_{m})}{\partial j_{2m-l}}=\sum_{i=0}^{\lfloor l/2\rfloor}{m-i\choose l-2i}\partial^{l-2i}_{-}B_{m-i}$ (4.4)
where $A_{n_{1}-i}$ and $B_{m-i}$ are functionals of $\ell_{n_{1}}$ and $\ell_{m}$, respectively. The binomial coefficients (${n\choose k}$ in the above expressions are understood to vanish when $k>n$. The explicit functional expression for $B_{m-i}$ is
@@
$+\sum_{k=1}^{i}\sum_{l=1}^{i-k}(-1)^{k+l}\left[{k+l-1\choose l}+2{k+l-2\choose l-1}\right]\partial_{-}^{k}\frac{\partial^{2}\ell_{m}}{\partial j_{m-i+k+l}\partial j_{m-i-l}}$ (4.5)
and likewise for $A_{n_{1}-i}$ after the replacements $m\to n_{1}$ and $\ell_{m}\to\ell_{n_{1}}$, with the understanding that in the above any derivative with respect to $j_{k}$ with $k<0$ is zero. The components $B_{m-i}$ of $\ell_{m}$ are such that the only non-mixed second order derivative present is, apart from $x^{-}$, with respect to $j_{m-i}$, and that $B_{m-i}=0$ if $\ell_{m}=\partial_{-}f_{m-1}$ is a total derivative of an order $m-1$ functional $f_{m-1}$, and likewise for $A_{n_{1}-i}$. We list here the first couple $A_{n_{1}-i}$ and $B_{m-i}$ in terms of $\ell_{n_{1}}$ and $\ell_{m}$
@@
$a^{q}_{ijk}$ =($={n_{1}-j\choose 2(i-j)-k}\Bigg{[}{2n_{1}+m-q-i+k+1\choose k+1}-{m-q+i\choose k+1}\Bigg{]}$ (4.8)
$b^{q}_{ijk}$ =($={m-q+i\choose 2(i-j)-k}\Bigg{[}{2(m-q)+n_{1}+j+k+1\choose k+1}-{n_{1}-j\choose k+1}\Bigg{]}$ (4.9)
@@
Let us try to understand the structure of this recursion relation in detail. In order for $B_{m-q}$, related to $\ell_{m}$ via (4), to be a local functional of $j(x^{-})$ and its derivatives, the right hand side of (4) multiplied by $A_{n_{1}}^{-\frac{2(m+n_{1}-q+1)}{2n_{1}+1}}$ must be a total derivative. Once this is the case, this equation can be freely integrated to solve for $B_{m-q}$, which is unique up to some integration constant. These integration constants, coming from solving for all $B_{m-i}$ up to $i=q$, amount to $\ell_{m}$ being written as a linear combination of functionals of lower order
@@
for which each ~$\tilde{\ell}_{m-i}$ commutes with $\ell_{n_{1}}$ independently of the others. Without loss of generality then, we may set $c_{0}=(-1)^{\frac{m-n_{1}}{2n_{1}+1}}$ and $c_{i}=0$ when $i\geq 1$ because a $c_{i}\neq 0$ amounts to solving for a functional $\ell_{m-i}$ of lower order which commutes with $\ell_{n_{1}}$. With this convention, $B_{m}$ can for instance be solved from (4) with $q=0$
@@
We find by directly solving for $B_{m-q}$ up to $q=4$ that the right hand side of (4) multiplied by $A_{n_{1}}^{-\frac{2(m+n_{1}-q+1)}{2n_{1}+1}}$ is in fact a total derivative, regardless of the values of $A_{n_{1}-i}$ with $0\leq i\leq q$, and therefore $\ell_{n_{1}}$. This is a nontrivial and important property of this recursion relation, depending sensitively on the coefficients $a^{q}_{ijk}$ and $b^{q}_{ijk}$, which at this stage we cannot prove for general $0\leq q\leq n_{1}$ for $n_{1}>4$, but we simply assume is true hereafter. We discuss more details of this property later in section 5. If this were not true, then this would imply a constraint on the $A_{n_{1}-i}$’s and therefore $\ell_{n_{1}}$ in order for the $B_{m-q}$’s to be local functionals of $j(x^{-})$ and its derivatives.
@@
Consider then the case when $\lfloor\frac{m-n_{1}-1}{3}\rfloor+1\geq n_{1}$, which is satisfied when $m\geq 4n_{1}-2$. In this case, $B_{m-q}$ is (4.13) for all $0\leq q<n_{1}$, and this fixes the $j_{m-q}$ dependence of $\ell_{m}$ by integrating (4.13) twice with respect to $j_{m-q}$
@@
where $f_{m-n_{1}}$ is some undetermined functional of order $m-n_{1}$. Here we have omitted in $\ell_{m}$ any term linear in $j_{m-q}$, because such a term is equivalent to a functional of order $m-q-1$.
@@
Note that this differs from (4) only by the addition of the term on the right hand side of the first line. If the recursion relation (4) satisfies the property that the right hand side multiplied by $A_{n_{1}}^{-\frac{2(m+n_{1}-q+1)}{2n_{1}+1}}$ is a total derivative, regardless of the values of $A_{n_{1}-i}$ with $0\leq i\leq q$ up to and including $q=n_{1}$, then consistency of (4) requires that the additional term multiplied by $A_{n_{1}}^{-\frac{2(m+1)}{2n_{1}+1}}$ is also a total derivative, i.e. after using (4.4), (4.12) and (4.6), the quantity
@@ 5 Discussion
In this paper, we found necessary conditions (3.12) on all possible infinite dimensional mutually commuting subalgebras of the $2d$ single scalar conformal field theory. In particular, we found the necessary $j_{n_{1}}$ dependence of the lowest order Lagrangian $\ell_{n_{1}}$ of any infinite dimensional mutually commuting subalgebra of (2.6) when $n_{1}=1,2,3,4$ and more generally for $n_{1}>4$ if the recursion relation (4) satisfies the property that the right hand side multiplied by $A_{n_{1}}^{-\frac{2(m+n_{1}-q+1)}{2n_{1}+1}}$ is a total derivative regardless of the values of all $A_{n_{1}-i}$, for all $0\leq q\leq n_{1}$.
@@
where the primes here indicate an $x^{-}$ derivative. For $q=2,3,4$, $B_{m-q}$ are more complicated, and so we provide them in the Mathematica ancillary file Bm-q.m. When $q>4$, the right hand side is sufficiently complicated so that it becomes too time consuming to check that this property is satisfied recursively.
@@
That the right hand side of (4) multiplied by $A_{n_{1}}^{-\frac{2(m+n_{1}-q+1)}{2n_{1}+1}}$, which we will call $D_{q}$ as a shorthand, is a total derivative then means that
[…]
Given that we already have a base case, proving that this is true for all $0\leq q\leq n_{1}$ by mathematical induction would amount to proving that if (5.4) is true for all $10\leq q\leq n-$ 1 given that $B_{m-j}$ satisfies the recursion relation (4) for all $10\leq j\leq n-$ 1, then it is also true at $q=n$. For example, checking this at $i=q=n$, we find
@@
which is zero by virtue of (4) at $q=0$. Because of the form of the recursion relation (4), checking that $\mathcal{E}^{n_{1}-i}(D_{q})=0$ at $i<q=n$ will require more and more input from (4) as $i$ gets smaller. In particular, checking that $\mathcal{E}^{n_{1}-i}(D_{q})=0$ requires knowledge of $B_{m-j}$ for all $0\leq j\leq q-i$ when $1\leq i\leq q$, and for all $10\leq j\leq q-$ 1 when $i=0$. This makes it difficult to prove by mathematical induction. Without a more promising direction for proving this property at this time, we will simply leave it here as a conjecture.
@@
Beyond these technical considerations, there are several open questions worth exploring in the future. For instance, while we have determined the necessary $j_{n_{1}}$ dependence of $\ell_{n_{1}}$, it will be interesting to determine what if any necessary constraints there are on the undetermined functionals $a_{n_{1}-1}(j,\dots,j_{n_{1}-1})$ and $b_{n_{1}-1}(j,\dots,j_{n_{1}-1})$ in (3.12). As we saw in Lindwasser:2024qyh , there are further constraints on $a_{n_{1}-1}$ and $b_{n_{1}-1}$ if one requires that $\ell_{n_{1}}$ commutes with $\ell_{m}$ of a given order $m$. However there may be other subalgebras with a sequence of Lagrangians of orders $n_{1},n_{2},\dots,n_{i},\dots$ with the same $n_{1}$ but all $n_{i}\neq m$. What we would like to find then are constraints on $a_{n_{1}-1}$ and $b_{n_{1}-1}$ which are independent of $m$, as we did for the $j_{n_{1}}$ dependence of $\ell_{n_{1}}$. To do so, it is necessary to impose (4.1) when $p>2n_{1}$.
@@
Given an infinite dimensional mutually commuting subalgebra with lowest nonzero order Lagrangian $\ell_{n_{1}}$, it would be interesting to determine what sequences $n_{1},n_{2},\dots,n_{i},\dots$ of orders of Lagrangians are allowed. Recall that when $n_{1}=1$, all subalgebras found obey $n_{i}\equiv 0,1\,(\text{mod 2})$, whereas when $n_{1}=2$, all subalgebras found obey $n_{i}\equiv 0,2\,(\text{mod 3})$. With the limited data we have, it appears there is a unique sequence of orders for each $n_{1}$. It is tempting to propose that this is true for all $n_{1}$, and that the sequence obeys $n_{i}\equiv 0,n_{1}\,(\text{mod }n_{2})$. Whether or not this is true would be interesting to pursue in the future. To begin with one should at least find subalgebras with $n_{1}>2$, perhaps by using the ansatz $\ell_{n_{1}}=(1+j_{n_{1}})^{\frac{1}{n_{1}+1}}$ to search for commuting Lagrangians with definite scaling dimension assuming $[j]=-n_{1}$ and $[x^{-}]=-1$. Interestingly, it was proven through other means that there are no infinite dimensional mutually commuting subalgebras of definite scaling dimension polynomial Lagrangians with $n_{1}>2$ when $[j]>0$ SANDERS1998410 . We would like to understand if this is true more generally in the future.
@@
The methods presented in this paper are best suited for narrowing in on the space of integrable deformations. They cannot at present prove that any given mutually commuting subalgebra is infinite dimensional. It may in fact be the case that a subalgebra defined by $\ell_{n_{1}}$ for a given $a_{n_{1}-1}$, $b_{n_{1}-1}$, and $c$ is either not infinite dimensional, or does not exist. Paired with these methods, we would like to develop methods of proving integrability of the models found in this paper.
@@
The same analysis can be done for any theory with left(right)-moving currents, such as a collection of scalars, a collection of left(right)-moving fermions, Wess–Zumino–Witten models, or principal chiral models. All of these theories have an infinite dimensional Lie algebra, whose infinite dimensional mutually commuting subalgebras define integrable hierarchies, which will be interesting to find. Finally, there is a wealth of literature on integrable deformations Klimcik:2008eq ; Klimcik:2014bta ; Sfetsos:2013wia ; Zarembo:2017muf ; Orlando:2019his ; Klimcik:2021bjy ; Hoare:2021dix ; Ferko:2024ali ; Bielli:2024fnp of these models which are not of the type we considered here, and it will be interesting to determine if the methods in this paper are related, or can perhaps be accommodated to reproduce, these integrable deformations.

### v3

@@ 4 $\mathcal{E}(\partial_{-}\mathcal{E}(\ell_{n_{1}})\mathcal{E}(\ell_{m}))=0$
To begin, recall that $\mathcal{E}(\partial_{-}\mathcal{E}(\ell_{n_{1}})\mathcal{E}(\ell_{m}))=0$ is a polynomial equation that involves $j(x^{-})$ and its derivatives up to order $2(n_{1}+m)$. So if $\mathcal{E}(\partial_{-}\mathcal{E}(\ell_{n_{1}})\mathcal{E}(\ell_{m}))=0$, then any derivative of $\mathcal{E}(\partial_{-}\mathcal{E}(\ell_{n_{1}})\mathcal{E}(\ell_{m}))$, say with respect to $j_{2(n_{1}+m)-p}$ with $0\leq p\leq 2(n_{1}+m)$, is zero
@@
We first focus on (4.1) when $0\leq p<2n_{1}$. Using the above and the commutation relation $\frac{\partial}{\partial j_{k}},\partial_{-}]=\frac{\partial}{\partial j_{k-1}}$, (4.1) becomes
@@
$\hskip-22.76228pt+\sum_{k=0}^{p}\sum_{r=0}^{k}\Bigg(\hskip-2.56064pt{2m-p+1+k\choose k+1-r}\partial_{-}^{k+1-r}\frac{\partial\mathcal{E}(\ell_{n_{1}})}{\partial j_{2n_{1}-r}}\frac{\partial\mathcal{E}(\ell_{m})}{\partial j_{2m-p+k}}-{2n_{1}+1-r\choose k+1-r}\partial_{-}^{k+1-r}\frac{\partial\mathcal{E}(\ell_{m})}{\partial j_{2m-p+k}}\frac{\partial\mathcal{E}(\ell_{n_{1}})}{\partial j_{2n_{1}-r}}\Bigg)$
@@
For $0\leq p<2n_{1}$, this is a recursion relation for $\frac{\partial\mathcal{E}(\ell_{m})}{\partial j_{2m-p}}$ of order $p$, which can be used to solve for $\frac{\partial\mathcal{E}(\ell_{m})}{\partial j_{2m-p}}$ in terms of $\frac{\partial\mathcal{E}(\ell_{n_{1}})}{\partial j_{2n_{1}-r}}$ with $0\leq r\leq p$.
We simplify the recursion relation further by expanding out the derivatives $\frac{\partial\mathcal{E}(\ell_{n_{1}})}{\partial j_{2n_{1}-r}}$ and $\frac{\partial\mathcal{E}(\ell_{m})}{\partial j_{2m-r}}$, and noting that they can be written in the following form
$\frac{\partial\mathcal{E}(\ell_{n_{1}})}{\partial j_{2n_{1}-r}}=\sum_{i=0}^{\lfloor r/2\rfloor}{n_{1}-i\choose r-2i}\partial^{r-2i}_{-}A_{n_{1}-i},$ $\frac{\partial\mathcal{E}(\ell_{m})}{\partial j_{2m-r}}=\sum_{i=0}^{\lfloor r/2\rfloor}{m-i\choose r-2i}\partial^{r-2i}_{-}B_{m-i}$ (4.4)
where $A_{n_{1}-i}$ and $B_{m-i}$ are functionals of $\ell_{n_{1}}$ and $\ell_{m}$, respectively. The binomial coefficients $n\choose k$} in the above expressions are understood to vanish when $k>n$. The explicit functional expression for $B_{m-i}$ is
@@
$+\sum_{k=1}^{i}\sum_{r=1}^{i-k}(-1)^{k+r}\left[{k+r-1\choose r}+2{k+r-2\choose r-1}\right]\partial_{-}^{k}\frac{\partial^{2}\ell_{m}}{\partial j_{m-i+k+r}\partial j_{m-i-r}}$ (4.5)
and likewise for $A_{n_{1}-i}$ after the replacements $m\to n_{1}$ and $\ell_{m}\to\ell_{n_{1}}$, with the understanding that in the above any derivative with respect to $j_{k}$ with $k<0$ is zero. The components $B_{m-i}$ of $\ell_{m}$ are such that the only non-mixed second order derivative present is, apart from $x^{-}$, with respect to $j_{m-i}$, and that $B_{m-i}=0$ if $\ell_{m}=\partial_{-}f_{m-1}$ is a total derivative of an order m1m-1 functional $f_{m-1}$, and likewise for $A_{n_{1}-i}$. We list here the first couple $A_{n_{1}-i}$ and $B_{m-i}$ in terms of $\ell_{n_{1}}$ and $\ell_{m}$,
@@
$a^{q}_{ijk}$ $={n_{1}-j\choose 2(i-j)-k}\Bigg[{2n_{1}+m-q-i+k+1\choose k+1}-{m-q+i\choose k+1}\Bigg$] (4.8)
$b^{q}_{ijk}$ $={m-q+i\choose 2(i-j)-k}\Bigg[{2(m-q)+n_{1}+j+k+1\choose k+1}-{n_{1}-j\choose k+1}\Bigg$] (4.9)
@@
Let us try to understand the structure of this recursion relation in detail. In order for $B_{m-q}$, related to $\ell_{m}$ via (4), to be a local functional of $j(x^{-})$ and its derivatives, the right-hand side of (4) multiplied by $A_{n_{1}}^{-\frac{2(m+n_{1}-q+1)}{2n_{1}+1}}$ must be a total derivative. Once this is the case, this equation can be freely integrated to solve for $B_{m-q}$, which is unique up to some integration constant. These integration constants, coming from solving for all $B_{m-i}$ up to $i=q$, amount to $\ell_{m}$ being written as a linear combination of functionals of lower order
@@
for which each $\tilde{\ell}_{m-i}$ commutes with $\ell_{n_{1}}$ independent of the others. Without loss of generality then, we may set $c_{0}=(-1)^{\frac{m-n_{1}}{2n_{1}+1}}$ and $c_{i}=0$ when $i\geq 1$ because a $c_{i}\neq 0$ amounts to solving for a functional $\ell_{m-i}$ of lower order that commutes with $\ell_{n_{1}}$. With this convention, $B_{m}$ can, for instance, be solved from (4) with $q=0$,
@@
We find by directly solving for $B_{m-q}$ up to $q=4$ that the right-hand side of (4) multiplied by $A_{n_{1}}^{-\frac{2(m+n_{1}-q+1)}{2n_{1}+1}}$ is, in fact, a total derivative, regardless of the values of $A_{n_{1}-i}$ with $0\leq i\leq q$ and therefore $\ell_{n_{1}}$. This is a nontrivial and important property of this recursion relation, depending sensitively on the coefficients $a^{q}_{ijk}$ and $b^{q}_{ijk}$, which at this stage we cannot prove for general $0\leq q\leq n_{1}$ for $n_{1}>4$, but we simply assume is true hereafter. We discuss more details of this property later in section 5. If this were not true, then this would imply a constraint on the $A_{n_{1}-i}$’s and therefore $\ell_{n_{1}}$ in order for the $B_{m-q}$’s to be local functionals of $j(x^{-})$ and its derivatives.
@@
Consider then the case when $\lfloor\frac{m-n_{1}-1}{3}\rfloor+1\geq n_{1}$, which is satisfied when $m\geq 4n_{1}-2$. In this case, $B_{m-q}$ is (4.13) for all $0\leq q<n_{1}$, and this fixes the $j_{m-q}$ dependence of $\ell_{m}$ by integrating (4.13) twice with respect to $j_{m-q}$,
@@
where $f_{m-n_{1}}$ is some undetermined functional of order $m-n_{1}$. Here we have omitted in $\ell_{m}$ any term linear in $j_{m-q}$, because such a term is equivalent to a functional of order mq1m-q-1.
@@
Note that this differs from (4) only by the addition of the term on the right-hand side of the first line. If the recursion relation (4) satisfies the property that the right-hand side multiplied by $A_{n_{1}}^{-\frac{2(m+n_{1}-q+1)}{2n_{1}+1}}$ is a total derivative, regardless of the values of $A_{n_{1}-i}$ with $0\leq i\leq q$ up to and including $q=n_{1}$, then consistency of (4) requires that the additional term multiplied by $A_{n_{1}}^{-\frac{2(m+1)}{2n_{1}+1}}$ is also a total derivative, i.e. after using (4.4), (4.12) and (4.6), the quantity
@@ 5 Discussion
In this paper, we found necessary conditions (3.12) on all possible infinite dimensional mutually commuting subalgebras of the $2d$ single scalar conformal field theory. In particular, we found the necessary $j_{n_{1}}$ dependence of the lowest order Lagrangian $\ell_{n_{1}}$ of any infinite dimensional mutually commuting subalgebra of (2.6) when $n_{1}=1,2,3,4$ and more generally for $n_{1}>4$ if the recursion relation (4) satisfies the property that the right-hand side multiplied by $A_{n_{1}}^{-\frac{2(m+n_{1}-q+1)}{2n_{1}+1}}$ is a total derivative regardless of the values of all $A_{n_{1}-i}$, for all $0\leq q\leq n_{1}$.
@@
where the primes here indicate an $x^{-}$ derivative. For $q=2,3,4$, $B_{m-q}$ are more complicated, and so we provide them in the Mathematica ancillary file Bm-q.m. When $q>4$, the right-hand side is sufficiently complicated so that it becomes too time consuming to check that this property is satisfied recursively.
@@
That the right-hand side of (4) multiplied by $A_{n_{1}}^{-\frac{2(m+n_{1}-q+1)}{2n_{1}+1}}$, which we will call $D_{q}$ as a shorthand, is a total derivative then means that
$\mathcal{E}^{n_{1}-i}(D_{q})=0,$  $\forall\,0\leq i\leq q$ (5.4)
Given that we already have a base case, proving that this is true for all $0\leq q\leq n_{1}$ by mathematical induction would amount to proving that if (5.4) is true for all $0\leq q\leq n-1$ given that $B_{m-j}$ satisfies the recursion relation (4) for all $0\leq j\leq n-1$, then it is also true at $q=n$. For example, checking this at $i=q=n$, we find
@@
which is zero by virtue of (4) at $q=0$. Because of the form of the recursion relation (4), checking that $\mathcal{E}^{n_{1}-i}(D_{q})=0$ at $i<q=n$ will require more and more input from (4) as $i$ gets smaller. In particular, checking that $\mathcal{E}^{n_{1}-i}(D_{q})=0$ requires knowledge of $B_{m-j}$ for all $0\leq j\leq q-i$ when $1\leq i\leq q$, and for all $0\leq j\leq q-1$ when $i=0$. This makes it difficult to prove by mathematical induction. Without a more promising direction for proving this property at this time, we will simply leave it here as a conjecture.
@@
Beyond these technical considerations, there are several open questions worth exploring in the future. For instance, while we have determined the necessary $j_{n_{1}}$ dependence of $\ell_{n_{1}}$, it will be interesting to determine what, if any, necessary constraints there are on the undetermined functionals $a_{n_{1}-1}(j,\dots,j_{n_{1}-1})$ and $b_{n_{1}-1}(j,\dots,j_{n_{1}-1})$ in (3.12). As we saw in Lindwasser:2024qyh , there are further constraints on $a_{n_{1}-1}$ and $b_{n_{1}-1}$ if one requires that $\ell_{n_{1}}$ commutes with $\ell_{m}$ of a given order $m$. However, there may be other subalgebras with a sequence of Lagrangians of orders $n_{1},n_{2},\dots,n_{i},\dots$ with the same $n_{1}$ but all $n_{i}\neq m$. What we would like to find then are constraints on $a_{n_{1}-1}$ and $b_{n_{1}-1}$ that are independent of $m$, as we did for the $j_{n_{1}}$ dependence of $\ell_{n_{1}}$. To do so, it is necessary to impose (4.1) when $p>2n_{1}$.
@@
Given an infinite dimensional mutually commuting subalgebra with lowest nonzero order Lagrangian $\ell_{n_{1}}$, it would be interesting to determine what sequences $n_{1},n_{2},\dots,n_{i},\dots$ of orders of Lagrangians are allowed. Recall that when $n_{1}=1$, all subalgebras found obey $n_{i}\equiv 0,1\,(\text{mod 2})$, whereas when $n_{1}=2$, all subalgebras found obey $n_{i}\equiv 0,2\,(\text{mod 3})$. With the limited data we have, it appears there is a unique sequence of orders for each $n_{1}$. It is tempting to propose that this is true for all $n_{1}$, and that the sequence obeys $n_{i}\equiv 0,n_{1}\,(\text{mod }n_{2})$. Whether or not this is true would be interesting to pursue in the future. To begin with one should at least find subalgebras with $n_{1}>2$, perhaps by using the ansatz $\ell_{n_{1}}=(1+j_{n_{1}})^{\frac{1}{n_{1}+1}}$ to search for commuting Lagrangians with definite scaling dimension assuming $j]=-n_{1}$ and $x^{-}]=-1$. Interestingly, it was proven through other means that there are no infinite dimensional mutually commuting subalgebras of definite scaling dimension polynomial Lagrangians with $n_{1}>2$ when $[j]>0$ SANDERS1998410 . We would like to understand if this is true more generally in the future.
@@
The methods presented in this paper are best suited for narrowing in on the space of integrable deformations. They cannot at present prove that any given mutually commuting subalgebra is infinite dimensional. It may, in fact, be the case that a subalgebra defined by $\ell_{n_{1}}$ for a given $a_{n_{1}-1}$, $b_{n_{1}-1}$, and $c$ is either not infinite dimensional, or does not exist. Paired with these methods, we would like to develop methods of proving integrability of the models found in this paper.
@@
The same analysis can be done for any theory with left- (right-) moving currents, such as a collection of scalars, a collection of left- (right-) moving fermions, Wess–Zumino–Witten models, or principal chiral models. All of these theories have an infinite dimensional Lie algebra, whose infinite dimensional mutually commuting subalgebras define integrable hierarchies, which will be interesting to find. Finally, there is a wealth of literature on integrable deformations Klimcik:2008eq ; Klimcik:2014bta ; Sfetsos:2013wia ; Zarembo:2017muf ; Orlando:2019his ; Klimcik:2021bjy ; Hoare:2021dix ; Ferko:2024ali ; Bielli:2024fnp of these models that are not of the type we considered here, and it will be interesting to determine if the methods in this paper are related, or can perhaps be accommodated to reproduce, these integrable deformations.